MuchQuestions111:

A marine biologist is studying the growth of a particular species of fish. She writes the following equation to show the length of the fish, f(m), in cm, after m months: f(m) = 4(1.08)m Part A: When the marine biologist concluded her study, the length of the fish was approximately 6.86 cm. What is a reasonable domain to plot the growth function? Part B: What does the y-intercept of the graph of the function f(m) represent? Part C: What is the average rate of change of the function f(m) from m = 3 to m = 7, and what does it represent?

Astrid1:

Hello and welcome. Asked and Answered here: https://questioncove.com/updates/5db1ab9d323a68b899c3f600

MuchQuestions111:

The question is different

MuchQuestions111:

This one has 1.08

MuchQuestions111:

The one you showed me has 1.09

MuchQuestions111:

in the question

Astrid1:

Okay then just follow the steps but do it with 1.08 instead.

MuchQuestions111:

-_- I've tried and I've been stuck for at least 37 mins

TETSXPREME:

Hello Welcome Do you still need help? If not please close

surjithayer:

\[6.86=4(1.08)^m,\]\[(1.08)^m=\frac{ 6.86 }{ 4}=1.715\] \[m \ln (.08)=\ln(1.715)\] \[m=\frac{ \ln 1.715 }{ \ln 1.08 }\approx7.0089\] so domain is 0<m<7.1

surjithayer:

b. when m=0 \[f(m)=4(1.08)^0=4\times1=4\] so y intercept is 4 average rate of change for m=3 to m=7 is \[=\frac{ f(7)-f(3) }{ 7-3 }\] \[=\frac{ 4(1.08)^7-4(1.08)^3 }{ 7-3 }\] \[=\frac{ 4(1.08)^3[(1.08)^4-1] }{ 4 }\] \[=(1.08)^3[1.08^4-1]\approx 0.454 cm\]